OpenStudy (anonymous):

help! I tried using the substitution method, it's not making sense. The problem says: What is the solution of the system of equations? 3x + 2y + z = 7 5x + 5y + 4z = 3 3x + 2y + 3z =1

OpenStudy (anonymous):

Ew. Just ew. Well, I would start with solving two parts. If you take 3x+2y+z=7 And subtract from that 3x+2y+3z=1 You get -2z=6 Now that you know z, can you find the rest?

OpenStudy (anonymous):

I figured out z by doing just what you did, but then I got confused from there cuz it wasnt making sense

OpenStudy (anonymous):

Well, you then get: 3x+2y=10 5x+5y=15 (I cleaned up the z's already) Solving for x in the first you can get x=(10-2y)/3 Put that into the second equation and find y. And go back and get x.

OpenStudy (anonymous):

oh ok. Thanks for your help!